Perform these operations with the appropriate number of significant figures. (3.517 ✕ 10^6) (4.8 ✕ 10^-2)
Answer #1

actually multiplication = 3.517* 4.8* 10^4 = 16.88 *10^4

Considering significant figures

=1.7*10^5

=======

Comment in case any doubt, will reply for sure Goodluck